# KCL-TH-2024-59

### v1

@@ I Introduction
Over the past 50 years, compact objects have acquired a center-stage role in many areas of astrophysics and fundamental physics. Neutron stars (NSs) Hewish et al. (1968); Shapiro and Teukolsky (1983) form the end product of the evolution of moderately massive stars and, through binary coalescence, give rise to short gamma-ray bursts Berger (2014); Abbott et al. (2017). Supermassive BHs residing at the center of most (if not all) galaxies are intricately related to their host dynamics and evolution Cattaneo et al. (2009) and can drive some of the most luminous objects in the universe Rees (1984). NSs and stellar-mass BHs, when accreting matter from a binary companion, manifest themselves as bright X-ray sources Reig (2011) and so-called primordial BHs are hypothesized as possible remnants of the early stages of cosmological evolution Escrivà et al. (2022). At a fundamental level, NSs represent unique laboratories to study the equation of state of matter at super-nuclear densities Özel and Freire (2016); Abbott et al. (2018); Lattimer (2021) while BHs, through their rotation rates, may provide us with information about matter fields beyond the standard model of particle physics Cardoso et al. (2015); Barack et al. (2019); Cardoso et al. (2018). Both, NSs and BHs enable us to test general relativity in the strong-field regime Berti et al. (2015); Abbott et al. (2021) and, as mathematical solutions, have been essential in deepening our understanding of Einstein’s theory of general relativity; the Schwarzschild solution in Kruskal-Szekeres coordinates or the Buchdahl limit are prominent points in case. On the observational side, the exploration of compact objects has received a tremendous boost from the direct detection of gravitational waves (GWs) by the Laser Interferometer Gravitational-Wave Observatory (LIGO) and Virgo Abbott et al. (2016, 2023).
@@
One of the most exciting questions in this context is the very nature of compact objects and whether these encompass exotic members composed of matter beyond the standard model of particle physics. Exotic compact objects may furthermore constitute a significant fraction of the enigmatic dark matter content of the universe, either themselves or as localized concentrations indicative of fundamental fields permeating spacetime on grander scales. While a plethora of experiments has been searching for dark-matter candidate fields Feng (2010); Kahlhoefer (2017); Cebrián (2023), their capacity to form gravitationally bound macroscopic concentrations or “stars” offers the unique opportunity to study them as compact-binary sources of GWs. Such an observational effort, however, requires a detailed theoretical understanding and precision modeling of exotic compact binaries for two main reasons. First, the canonical approach to analyze observational data from GW detectors is based on Bayesian statistics and matched filtering techniques employing extensive theoretical GW template banks Jaranowski and Krolak (2005); Maggiore (2007). Second, as evidenced in Ref. Evstafyeva et al. (2024) for the case of boson stars and the sensitivity of present GW detectors, the GW signals generated by the inspiral and coalescence of exotic compact binaries can exhibit significant degeneracy with those from black-hole (BH) binaries. Breaking this degeneracy through the identification of characteristic signatures from exotic compact objects will be of great benefit in our GW guided quest for new physics.
@@
The main goal of this paper is to explore in depth specific characteristics of the GW emission from stars composed of a complex scalar field. This class of compact objects, henceforth referred to simply as boson stars (BSs) (see e.g. Refs. Liebling and Palenzuela (2012); Visinelli (2021); Bezares and Sanchis-Gual (2024) for reviews), combines a deceptive simplicity regarding their mathematical foundation with a remarkably rich phenomenology of their dynamics and radiative properties. As such, BSs are not only interesting in their own right, but also provide an excellent proxy to capture main effects for a wider range of compact objects. Put another way, should the analysis of a future GW event exhibit statistically significant preference for BS templates relative to BH ones, the impact would be substantial whether or not the source eventually turns out to be a BS or another exotic source as for example cosmic strings Aurrekoetxea et al. (2024) or compact stars composed of higher-spin fields Calderón Bustillo et al. (2021).
@@
The concept of BSs first appeared almost 60 years ago thanks to the pioneering work by Kaup Kaup (1968), whilst their dynamics in binary systems have been studied numerically with increasing intensity only for about 20 years. These studies have already uncovered a considerable amount of valuable insight on the properties of these exotic objects by focusing on binaries consisting of rotating stars Siemonsen and East (2021); Sanchis-Gual et al. (2019a); Yoshida and Eriguchi (1997); Mielke (2016); Kleihaus et al. (2005); Di Giovanni et al. (2020); Dmitriev et al. (2021); Cardoso et al. (2008); Lai (2004), BSs with light rings Siemonsen (2024); Cunha et al. (2017); Cardoso et al. (2016); Cunha et al. (2023), equal- Helfer et al. (2022); Croft et al. (2023); Palenzuela et al. (2017); Atteneder et al. (2024); Bezares et al. (2017); Sanchis-Gual et al. (2020); Palenzuela et al. (2008); Mundim (2010) and unequal-mass binaries Evstafyeva et al. (2023a); Siemonsen and East (2023a); Bezares et al. (2022), hybrid BS-NS Dietrich et al. (2019) and BS-BH Clough et al. (2018); Cardoso et al. (2022) binaries, multiple scalar fields Bezares and Palenzuela (2018); Alcubierre et al. (2018); Sanchis-Gual et al. (2021); Jaramillo et al. (2020); Bernal et al. (2010); Hawley and Choptuik (2003); Guzmán and Avilez (2018), real scalar fields (aka oscillatons) Helfer et al. (2019); Brito et al. (2016a); Widdicombe et al. (2020); Grandclement et al. (2011) and vector fields (aka Proca stars) Brito et al. (2016b); Sanchis-Gual et al. (2019b); Calderón Bustillo et al. (2021); Herdeiro et al. (2024). The BS parameter space offers ample avenues for further exploration. For example, the impact of dephasing and the vortex structure on the dynamics of inspiralling BS binaries has only recently been studied in Ref. Siemonsen and East (2023b). Their results demonstrate for the first time that a rotating BS can form from the coalescence of two non-spinning BSs. Formation of these objects are less well studied, but in Widdicombe et al. (2018), it was shown that real scalar BS can generically form from the collapse of overdense regions. Similarly, thanks to increased efforts to model exotic compact objects more extensively, a surrogate model for equal-mass head-on collisions of Proca stars has been constructed in Ref. Luna et al. (2024).
@@
The merger dynamics of BS binaries have been found to exhibit a (perhaps surprisingly) rich phenomenology driven by the specific properties of the individual BSs. For instance, Ref. Palenzuela et al. (2007) observes that the GW emission of the simulated head-on collisions of equal-mass mini BSs is quadrupole dominated but with varying characteristics. On the one hand, in-phase BSs merge into a single more massive star whereas, on the other hand, binaries with maximally dephased scalar fields come close to each other but remain separate, almost like two touching snooker balls. For large initial momenta, the two mini BSs can even pass through each other Choi et al. (2009); Widdicombe et al. (2020), exhibiting a true solitonic nature. An even more exotic fate of an inspiralling BS binary has been found in Ref. Siemonsen and East (2023a), where an aligned-spin binary, instead of merging, forms a binary of nonspinning BSs flung out away from the center of mass. Ultrarelativistic BS collisions were modeled in Ref. Choptuik and Pretorius (2010) and found to result in BH formation, consistent with predictions by Thorne’s hoop conjecture Thorne (1972).
@@
Numerical explorations of BSs have thus demonstrated many interesting features of their mergers and dynamics; however, our insight is largely limited to isolated points or patches in the BS parameter space leaving systematic explorations of the complete parameters as a key challenge for complementary work. The most comprehensive study in this direction has, to our awareness, been performed for Proca stars in Ref. Sanchis-Gual et al. (2022), covering a range of frequency (i.e. mass) ratios and phase offsets. The central goal of our work is to extend the parameter coverage to the potential function for head-on collisions of equal-mass non-spinning scalar BSs. For this purpose we simulate $\sim 10^{4}$ BS merger systems composed of a scalar field of mass $m$ self-interacting through a solitonic potential
@@ II Executive summary
Figure 1: Left: The GW energy of the head-on collision of two BSs with a solitonic potential (1) for $\sigma=0.27825M_{\rm{Pl}}$ is shown as a function of the BSs’ central scalar-field amplitude $|\phi_{c}|$. Blue dots represent binaries completing their infall as BSs and forming a BH or BS at merger whereas red symbols denote systems where each BS collapses to a BH prior to merger. The purple line represents the mass curve $M(|\phi_{c}|)$ for single spherical BSs with central scalar-field amplitude $|\phi_{c}|$. Right: Zoom into the region around the sharp variations of $E_{\rm GW}(|\phi_{\rm c}|)$.
@@
    BS collisions can be radiatively way more efficient than those of BHs. The GW energy released in the BS collisions reaches values up to $E_{\rm GW}/M_{\rm tot}=0.0035$, about six times larger than the corresponding number $6\times 10^{-4}$ observed in head-on collisions of non-spinning equal-mass BHs; see e.g. Fig. 4 in Ref. Sperhake et al. (2020) and also compare with the red dots in our Fig. 1 where the BSs have collapsed to BHs prior to merger. We will encounter yet larger energies when we discuss other $\sigma$ values later on. This effect was first noted in Ref. Helfer et al. (2019) in the context of self-gravitating real-scalar-field oscillatons, suggesting that this is a generic effect for mergers of compact objects composed of fundamental fields.
@@
    Even away from these discontinuities, small differences in the initial configuration lead to large, albeit continuous, ${\cal O}(1)$ differences in the GW emission. The functional dependence of the GW energy on $|\phi_{c}|$ exhibits sharp local extrema, most notably the local minimum at $|\phi_{\rm c}|/M_{\rm Pl}\approx 0.115$; we will find these sharp extrema – maxima or minima – to be even more prominent for other $\sigma$ values and, henceforth, refer to them as needles. This observation suggests that it can be misleading to assume similar BS collisions produce approximately equal output in GWs.
@@ III Theory and computational framework
where $|\phi(r)|$ is the scalar-field amplitude and $\omega$ the (constant positive) frequency. This ansatz can be generalized by writing the phase factor as $e^{\mathrm{i}\mkern 1.0mu(\epsilon\omega t+\delta\vartheta)}$, where $\epsilon=\pm 1$ distinguishes BSs and anti-BSs, and $\delta\vartheta$ denotes a phase shift. In this work, we exclusively consider BS binaries without phase offset and, hence, set $\epsilon=1$, $\delta\vartheta=0$. With the decomposition (4), the Einstein equations for spherically symmetric BSs result in a coupled set of four first-order ordinary differential equations which we solve with a shooting method as detailed in Sec. 2.3 of Ref. Helfer et al. (2022); cf. also Appendix A of Ref. Evstafyeva et al. (2023b) for an in-depth discussion of the solution’s asymptotic behaviour at infinity.
For BS binary evolutions, we construct initial data by combining two such single-BS spacetimes using the constant volume element technique of Refs. Helfer et al. (2019, 2022); Evstafyeva et al. (2023a); cf. also Ref. Atteneder et al. (2024). The resulting binaries are characterized by the parameter $\sigma$ in the solitonic potential (1), the stars’ central scalar-field amplitudes $|\phi_{c}|$, their initial separation $D$ and their velocities $v$. Our study focuses on non-spinning equal-mass BS binaries, so that both stars have equal $|\phi_{c}|$ and, in the rest frame used for all our runs, their initial velocity has equal magnitude with direction pointing toward each other. Unless specified otherwise, we fix $v=0.1$ and $D=80\,m^{-1}$ which leaves us with two control parameters, $\sigma$ and $|\phi_{c}|$. In this paper, we present results for $\sigma/M_{\rm Pl}=\{0.25,0.28725,0.3,0.5\}$ as well as the mini-BS limit $\sigma\rightarrow\infty$; these cases are fully representative of the dynamics under consideration, but we note that further values of $\sigma$ are discussed in Ref.Ge (2024).
@@
We simulate the head-on collisions resulting from these initial data using the grchombo code Clough et al. (2015); Radia et al. (2022); Andrade et al. (2021) which evolves the Einstein equations in the CCZ4 formulation Alic et al. (2012). Our specific version of the CCZ4 equations is given in Sec. 2 of Ref. Radia et al. (2022) and is discretized using the method of lines with fourth-order finite differencing in space and a fourth-order Runge-Kutta scheme in time. grchombo implements Berger-Rigoutsos adaptive-mesh refinement provided by the chombo library Adams et al. (2019) and facilitates various tagging criteria for the grid generation as summarized in Sec. 3 of Ref. Radia et al. (2022). In this paper, we utilize the scalar field $|\phi|$ and the conformal factor $\chi$ for regridding with their respective thresholds set to 0.3. We exploit the inherent axisymmetry of our BS binary configurations by reducing our computational domain from three to two spatial dimensions using the modified cartoon method in the form detailed in Ref. Cook et al. (2016); cf. also Alcubierre et al. (2001); Pretorius (2005) as well as Appendix A for our cartoon implementation of the wave extraction. This combination of mesh refinement and dimensional reduction enables us to simulate approximately 10 000 simulations – each utilizing two nodes (56 logical cores) and lasting approximately four hours. We employ a domain of length $512\,m^{-1}$ and a grid setup consisting of a nested set of $L=7$ refinement levels with grid spacing $\mathrm{d}x=1/(32\,m)$ on the innermost level, increasing by a factor 2 consecutively on each level further out.
For the calibration of our numerical uncertainty, we have simulated a representative BS binary configuration with $\sigma=0.25\,M_{\rm Pl}$ and $|\phi_{\mathrm{c}}|=0.18\,M_{\rm Pl}$ using three resolutions, our default $N=256$ as well as two higher resolutions $N=512$ and $N=1024$. The result, displayed in Fig. 2, demonstrates convergence between second and third order. We conservatively estimate the discretization error by assuming second-order convergence for the Richardson extrapolation and obtain a relative error of about $2.5\,\%$ for $N=256$. By fitting the energy values obtained at different extraction radii at order $1/R_{\rm ex}$, we find an additional uncertainty of $\sim 2\,\%$ at our fiducial $R_{\rm ex}=140\,m^{-1}$ resulting in a total error budget of $\sim 4.5\,\%$.
@@
As a final test, we compare in Fig. 3 the GW signal for a BS head-on collision for $\sigma=0.35\,M_{\rm Pl}$ and $|\phi_{\rm c}|=0.086\,M_{\rm Pl}$ calculated with our modified Cartoon approach with the result from a fully three-dimensional computation. The results agree within about a percent, commensurate with the uncertainty of the three-dimensional simulation Helfer et al. (2022).

### v2

@@ I Introduction
Over the past 50 years, compact objects have acquired a center-stage role in many areas of astrophysics and fundamental physics. Neutron stars (NSs) Hewish:1968bj; Shapiro1983 form the end product of the evolution of moderately massive stars and, through binary coalescence, give rise to short gamma-ray bursts Berger:2013jza; LIGOScientific:2017ync. Supermassive BHs residing at the center of most (if not all) galaxies are intricately related to their host dynamics and evolution Cattaneo:2009ub and can drive some of the most luminous objects in the universe Rees:1984si. NSs and stellar-mass BHs, when accreting matter from a binary companion, manifest themselves as bright X-ray sources Reig:2011zp and so-called primordial BHs are hypothesized as possible remnants of the early stages of cosmological evolution Escriva:2022duf. At a fundamental level, NSs represent unique laboratories to study the equation of state of matter at super-nuclear densities Ozel:2016oaf; LIGOScientific:2018cki; Lattimer:2021emm while BHs, through their rotation rates, may provide us with information about matter fields beyond the standard model of particle physics Cardoso:2014uka; Barack:2018yly; Cardoso:2018tly. Both, NSs and BHs enable us to test general relativity in the strong-field regime Berti:2015itd; LIGOScientific:2021sio and, as mathematical solutions, have been essential in deepening our understanding of Einstein’s theory of general relativity; the Schwarzschild solution in Kruskal-Szekeres coordinates or the Buchdahl limit are prominent points in case. On the observational side, the exploration of compact objects has received a tremendous boost from the direct detection of gravitational waves (GWs) by the Laser Interferometer Gravitational-Wave Observatory (LIGO) and Virgo Abbott:2016blz; KAGRA:2021vkt.
@@
One of the most exciting questions in this context is the very nature of compact objects and whether these encompass exotic members composed of matter beyond the standard model of particle physics. Exotic compact objects may furthermore constitute a significant fraction of the enigmatic dark matter content of the universe, either themselves or as localized concentrations indicative of fundamental fields permeating spacetime on grander scales. While a plethora of experiments has been searching for dark-matter candidate fields Feng:2010gw; Kahlhoefer:2017dnp; Cebrian:2022brv, their capacity to form gravitationally bound macroscopic concentrations or “stars” offers the unique opportunity to study them as compact-binary sources of GWs. Such an observational effort, however, requires a detailed theoretical understanding and precision modeling of exotic compact binaries for two main reasons. First, the canonical approach to analyze observational data from GW detectors is based on Bayesian statistics and matched filtering techniques employing extensive theoretical GW template banks Jaranowski:2005hz; Maggiore:2007ulw. Second, as evidenced in Ref. Evstafyeva:2024qvp for the case of boson stars and the sensitivity of present GW detectors, the GW signals generated by the inspiral and coalescence of exotic compact binaries can exhibit significant degeneracy with those from black-hole (BH) binaries. Breaking this degeneracy through the identification of characteristic signatures from exotic compact objects will be of great benefit in our GW guided quest for new physics.
@@
The main goal of this paper is to explore in depth specific characteristics of the GW emission from stars composed of a complex scalar field. This class of compact objects, henceforth referred to simply as boson stars (BSs) (see e.g. Refs. Liebling:2012fv; Visinelli:2021uve; Bezares:2024btu for reviews), combines a deceptive simplicity regarding their mathematical foundation with a remarkably rich phenomenology of their dynamics and radiative properties. As such, BSs are not only interesting in their own right, but also provide an excellent proxy to capture main effects for a wider range of compact objects. Put another way, should the analysis of a future GW event exhibit statistically significant preference for BS templates relative to BH ones, the impact would be substantial whether or not the source eventually turns out to be a BS or another exotic source as for example cosmic strings Aurrekoetxea:2023vtp or compact stars composed of higher-spin fields CalderonBustillo:2020fyi.
@@
The concept of BSs first appeared almost 60 years ago thanks to the pioneering work by Kaup Kaup:1968zz, whilst their dynamics in binary systems have been studied numerically with increasing intensity only for about 20 years. These studies have already uncovered a considerable amount of valuable insight on the properties of these exotic objects by focusing on binaries consisting of rotating stars Siemonsen:2020hcg; Sanchis-Gual:2019ljs; Yoshida:1997qf; Mielke:2016war; Kleihaus:2005me; DiGiovanni:2020ror; Dmitriev:2021utv; Cardoso:2007az; Lai:2004fw, BSs with light rings Siemonsen:2024snb; Cunha:2017qtt; Cardoso:2016rao; Cunha:2022gde, equal- Helfer:2021brt; Croft:2022bxq; Palenzuela:2017kcg; Atteneder:2023pge; Bezares:2017mzk; Sanchis-Gual:2020mzb; Palenzuela:2007dm; Mundim:2010hi and unequal-mass binaries Evstafyeva:2022bpr; Siemonsen:2023age; Bezares:2022obu, hybrid BS-NS Dietrich:2018bvi and BS-BH Clough:2018exo; Cardoso:2022vpj binaries, multiple scalar fields Bezares:2018qwa; Alcubierre:2018ahf; Sanchis-Gual:2021edp; Jaramillo:2020rsv; Bernal:2009zy; Hawley:2002zn; Guzman:2018evm, real scalar fields (aka oscillatons) Helfer:2018vtq; Brito:2015yfh; Widdicombe:2019woy; Grandclement:2011wz and vector fields (aka Proca stars) Brito:2015pxa; Sanchis-Gual:2018oui; CalderonBustillo:2020fyi; Herdeiro:2023wqf. The BS parameter space offers ample avenues for further exploration. For example, the impact of dephasing and the vortex structure on the dynamics of inspiralling BS binaries has only recently been studied in Ref. Siemonsen:2023hko. Their results demonstrate for the first time that a rotating BS can form from the coalescence of two non-spinning BSs. Formation of these objects are less well studied, but in Widdicombe:2018oeo, it was shown that real-scalar stars can generically form from the collapse of overdense regions. Similarly, thanks to increased efforts to model exotic compact objects more extensively, a surrogate model for equal-mass head-on collisions of Proca stars has been constructed in Ref. Luna:2024kof.
@@
The merger dynamics of BS binaries have been found to exhibit a (perhaps surprisingly) rich phenomenology driven by the specific properties of the individual BSs. For instance, Ref. Palenzuela:2006wp observes that the GW emission of the simulated head-on collisions of equal-mass mini BSs is quadrupole dominated but with varying characteristics. On the one hand, in-phase BSs merge into a single more massive star whereas, on the other hand, binaries with maximally dephased scalar fields come close to each other but remain separate, almost like two touching snooker balls. For large initial momenta, the two mini BSs can even pass through each other Choi:2009; Widdicombe:2019woy, exhibiting a true solitonic nature. An even more exotic fate of an inspiralling BS binary has been found in Ref. Siemonsen:2023age, where an aligned-spin binary, instead of merging, forms a binary of nonspinning BSs flung out away from the center of mass. Ultrarelativistic BS collisions were modeled in Ref. Choptuik:2009ww and found to result in BH formation, consistent with predictions by Thorne’s hoop conjecture Thorne:1972ji.
@@
Numerical explorations of BSs have thus demonstrated many interesting features of their mergers and dynamics; however, our insight is largely limited to isolated points or patches in the BS parameter space leaving systematic explorations of the complete parameters as a key challenge for complementary work. The most comprehensive study in this direction has, to our awareness, been performed for Proca stars in Ref. Sanchis-Gual:2022mkk, covering a range of frequency (i.e. mass) ratios and phase offsets. The central goal of our work is to extend the parameter coverage to the potential function for head-on collisions of equal-mass non-spinning scalar BSs. For this purpose we simulate $\sim 10^{4}$ BS merger systems composed of a scalar field of mass $m$ self-interacting through a solitonic potential
@@ II Executive summary
Figure 1: Left: The GW energy of the head-on collision of two BSs with a solitonic potential (1) for $\sigma=0.27825M_{\rm{Pl}}$ is shown as a function of the BSs’ central scalar-field amplitude $|\phi_{\rm c}|$. Blue dots represent binaries completing their infall as BSs and forming a BH or BS at merger whereas red symbols denote systems where each BS collapses to a BH prior to merger. The numbers adjacent to the dots give the final digit of their $|\phi_{\rm c}|$ value. The purple line represents the mass curve $M(|\phi_{\rm c}|)$ for single spherical BSs with central scalar-field amplitude $|\phi_{\rm c}|$. Right: Zoom into the region around the sharp variations of $E_{\rm GW}(|\phi_{\rm c}|)$.
@@
    BS collisions can be radiatively way more efficient than those of BHs. The GW energy released in the BS collisions reaches values up to $E_{\rm GW}/M_{\rm tot}=0.0035$, about six times larger than the corresponding number $6\times 10^{-4}$ observed in head-on collisions of non-spinning equal-mass BHs; see e.g. Fig. 4 in Ref. Sperhake:2019oaw and also compare with the red dots in our Fig. 1 where the BSs have collapsed to BHs prior to merger. We will encounter yet larger energies when we discuss other $\sigma$ values later on. This effect was first noted in Ref. Helfer:2018vtq in the context of self-gravitating real-scalar-field oscillatons, suggesting that this is a generic effect for mergers of compact objects composed of fundamental fields.
@@
    Even away from these discontinuities, small differences in the initial configuration lead to large, albeit continuous, ${\cal O}(1)$ differences in the GW emission. The functional dependence of the GW energy on $|\phi_{\rm c}|$ exhibits sharp local extrema, most notably the local minimum at $|\phi_{\rm c}|/M_{\rm Pl}\approx 0.115$; we will find these sharp extrema – maxima or minima – to be even more prominent for other $\sigma$ values and, henceforth, refer to them as needles. This observation suggests that it can be misleading to assume similar BS collisions produce approximately equal output in GWs.
@@ III Theory and computational framework
where $|\phi(r)|$ is the scalar-field amplitude and $\omega$ the (constant positive) frequency. This ansatz can be generalized by writing the phase factor as $e^{\mathrm{i}\mkern 1.0mu(\epsilon\omega t+\delta\vartheta)}$, where $\epsilon=\pm 1$ distinguishes BSs and anti-BSs, and $\delta\vartheta$ denotes a phase shift. In this work, we exclusively consider BS binaries without phase offset and, hence, set $\epsilon=1$, $\delta\vartheta=0$. With the decomposition (4), the Einstein-Klein-Gordon equations for a spherically symmetric metric
$\mathrm{d}s^{2}=-\alpha(r)^{2}\mathrm{d}t^{2}+X(r)^{2}\mathrm{d}r^{2}+r^{2}(\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\varphi^{2})\,$,
result in a coupled set of four first-order ordinary differential equations which we solve with a shooting method as detailed in Sec. 2.3 of Ref. Helfer:2021brt; cf. also Appendix A of Ref. Evstafyeva:2023kfg for an in-depth discussion of the solution’s asymptotic behaviour at infinity.
For BS binary evolutions, we construct initial data by combining two such single-BS spacetimes using the constant volume element technique of Refs. Helfer:2018vtq; Helfer:2021brt; Evstafyeva:2022bpr; cf. also Ref. Atteneder:2023pge. This technique results in a significant improvement in the Hamiltonian constraint equation over the simple linear superposition of the two BS solutions.
The resulting binaries are characterized by the parameter $\sigma$ in the solitonic potential (1), the stars’ central scalar-field amplitudes $|\phi_{\rm c}|$, their initial separation $D$ and their velocities $v$. Our study focuses on non-spinning equal-mass BS binaries, so that both stars have equal $|\phi_{\rm c}|$ and, in the rest frame used for all our runs, their initial velocity has equal magnitude with direction pointing toward each other. Unless specified otherwise, we fix $v=0.1$ and $D=80\,m^{-1}$ which leaves us with two control parameters, $\sigma$ and $|\phi_{\rm c}|$. In this paper, we present results for $\sigma/M_{\rm Pl}=\{0.25,0.28725,0.3,0.5\}$ as well as the mini-BS limit $\sigma\rightarrow\infty$; these cases are fully representative of the dynamics under consideration, but we note that further values of $\sigma$ are discussed in Ref.Ge2024.
@@
We simulate the head-on collisions resulting from these initial data using the grchombo code Clough:2015sqa; Radia:2021smk; Andrade:2021rbd which evolves the Einstein equations in the CCZ4 formulation Alic:2011gg. Our specific version of the CCZ4 equations is given in Sec. 2 of Ref. Radia:2021smk and is discretized using the method of lines with fourth-order finite differencing in space and a fourth-order Runge-Kutta scheme in time. grchombo implements Berger-Rigoutsos adaptive-mesh refinement provided by the chombo library chombo and facilitates various tagging criteria for the grid generation as summarized in Sec. 3 of Ref. Radia:2021smk. In this paper, we utilize the scalar field $|\phi|$ and the conformal factor $\chi$ for regridding with their respective thresholds set to 0.3. We exploit the inherent axisymmetry of our BS binary configurations by reducing our computational domain from three to two spatial dimensions using the modified cartoon method in the form detailed in Ref. Cook:2016soy; cf. also Alcubierre:1999ab; Pretorius:2004jg as well as Appendix B for our cartoon implementation of the wave extraction. This combination of mesh refinement and dimensional reduction enables us to simulate approximately 10 000 simulations – each utilizing two nodes (56 logical cores) and lasting approximately four hours. We employ a domain of length $512\,m^{-1}$ and a grid setup consisting of a nested set of $L=7$ refinement levels with grid spacing $\mathrm{d}x=1/(32\,m)$ on the innermost level, increasing by a factor 2 consecutively on each level further out.
For the calibration of our numerical uncertainty, we have simulated a representative BS binary configuration with $\sigma=0.25\,M_{\rm Pl}$ and $|\phi_{\mathrm{c}}|=0.18\,M_{\rm Pl}$ using three resolutions, our default $N=256$ as well as two higher resolutions $N=512$ and $N=1024$. The result, displayed in Fig. 2, demonstrates convergence between second and third order. We conservatively estimate the discretization error by assuming second-order convergence for the Richardson extrapolation and obtain a relative error of about $2.5\,\%$ for $N=256$. By fitting the energy values obtained at different extraction radii with a polynomial in $1/R_{\rm ex}$ truncated at second order, we find an additional uncertainty of $\sim 2\,\%$ at our fiducial $R_{\rm ex}=140\,m^{-1}$ resulting in a total error budget of $\sim 4.5\,\%$. We have also monitored the constraint violations of our simulations as described in Appendix A.
@@
As a final test, we compare in Fig. 3 the GW signal for a BS head-on collision for $\sigma=0.35\,M_{\rm Pl}$ and $|\phi_{\rm c}|=0.086\,M_{\rm Pl}$ calculated with our modified Cartoon approach with the result from a fully three-dimensional computation. The results agree within about a percent, commensurate with the uncertainty of the three-dimensional simulation Helfer:2021brt.